OpenStudy (anonymous):

How many combinations of 5 students can a teacher choose from 30 students?

OpenStudy (shelby1290):

the teacher can choose 6 combinations of 5 students

OpenStudy (shelby1290):

30 divided by 5 = 6

OpenStudy (shelby1290):

@Venomous^^^ and to double check do 6x5 and you'll get 30 so 6 combinations of 5 students is correct!

OpenStudy (anonymous):

it is a permutation question, and i think it is also a combination question.

OpenStudy (anonymous):

How many combinations of 5 students can a teacher choose from 30 students? (Points : 4) 150 142,506 810,000 17,100,720

OpenStudy (anonymous):

What is the value of 7P3 ? (Points : 4) 35 56 210 840

OpenStudy (anonymous):

and i need these answers fast so please help me!!! thank you

OpenStudy (shelby1290):

for your first question the answer is 142,506

OpenStudy (shelby1290):

and for your second question it's 210 because... 7P3= 7!/(7-3)! = 7!/4! = 7*6*5 = 210

OpenStudy (shelby1290):

@Venomous there^

OpenStudy (anonymous):

WOW! Thank you @shelby1290 , you really saved me!!! thanks again!!! :)

OpenStudy (anonymous):

Felipe can choose 1 snack from a box containing 3 different granola bars, 4 different kinds of fruit, and 2 different kinds of chips. How many ways can Felipe choose a snack? (Points : 4) 5 9 14 24 Can you help me with this one to real quick? thanks

OpenStudy (shelby1290):

i think its 24

OpenStudy (anonymous):

okay thank you!